OpenStudy (anonymous):

What is the 20th term of the arithmetic sequence 3, –3, –9, … ?

OpenStudy (dan815):

whats the pattern

OpenStudy (dan815):

whats the 4th number tell me

OpenStudy (anonymous):

-15?

OpenStudy (dan815):

ok good so its losing 6 everytime

OpenStudy (dan815):

can you write an equation that is losing 6 every number and starts from 3

OpenStudy (anonymous):

20 = 9 - 6*20

OpenStudy (dan815):

hmm

OpenStudy (anonymous):

so it would be -111?

OpenStudy (dan815):

yep works

OpenStudy (anonymous):

k thanks

OpenStudy (dan815):

good job

OpenStudy (dan815):

just a suggest for future

OpenStudy (dan815):

write it like Sn = 9-6n Sn meaning number is sequence

OpenStudy (dan815):

when you write it in terms of n you can do quick checks to see if the equation works

OpenStudy (anonymous):

ya, i put 20 for the sn

OpenStudy (dan815):

S1 = 9-6(1) = 3 works for 1st term and so on

OpenStudy (dan815):

im just try to get you to start with these habits, it wont seem like much to you, but ull see its gonna clear up your work a lot, and ull start to think more like a mathematician

OpenStudy (anonymous):

thanks so much

OpenStudy (dan815):

anytime :)